# Learning Force Distribution Estimation for the GelSight Mini Optical Tactile Sensor Based on Finite Element Analysis

### v1

@@ V Conclusion
We introduced FEATS—a machine learning approach for estimating force distributions using the GelSight Mini tactile sensor. By training a U-net model on FEA-derived data, we achieved accurate predictions of both shear and normal forces from raw sensor images, with Mean Absolute Error (MAE) under $0.4\,\mathrm{N}$ on average in the range $0-40\,\mathrm{N}$ for the total normal force and $-5\,\mathrm{N}$ to $5\,\mathrm{N}$ for total shear force in each direction. The model shows potential for real-time applications and generalization to different sensors, although additional improvements are needed to account for the domain gap when a different sensor is used, which could be addressed, e.g., via the recently proposed domain adaptation for optical tactile sensors [46]. Our FEATS method offers an efficient online approach that amortizes the cost of running inverse FEA into a forward pass of a neural network, resulting in physically-grounded, interpretable representations for optical tactile sensors. Future work will focus on enhancing the model’s robustness to different tactile scenarios, refining its generalization capabilities across a broader range of objects and sensor types, and work towards demonstrating the representation’s effectiveness in robotic manipulation tasks.

### v2

@@ V Conclusion
We introduced FEATS—a machine learning approach for estimating force distributions using the GelSight Mini tactile sensor. By training a U-net model on FEA-derived data, we achieved accurate predictions of both shear and normal forces from raw sensor images, with Mean Absolute Error (MAE) under $0.8\,\mathrm{N}$ on average in the range $0-40\,\mathrm{N}$ for the total normal force and under $0.3\,\mathrm{N}$ on average for the total shear force in each direction with a range of $-5\,\mathrm{N}$ to $5\,\mathrm{N}$. The model shows potential for real-time applications and generalization to different indenters and sensors, although additional improvements are needed to account for the domain gap when a different sensor is used. This could be addressed, e.g., via the recently proposed domain adaptation for optical tactile sensors [49]. Our FEATS method offers an efficient online approach that amortizes the cost of running inverse FEA into a forward pass of a neural network, resulting in physically-grounded, interpretable representations for optical tactile sensors. Future work will focus on enhancing the model’s robustness to different tactile scenarios, refining its generalization capabilities across a broader range of objects and sensor types, and work towards demonstrating the representation’s effectiveness in robotic manipulation tasks.